# Causal Concept Embedding Models: Beyond Causal Opacity in Deep Learning

### v1

@@ 3.1 Blueprint
The distribution $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i})$ can be associated with a structural causal model =($\mathcal{M}=(\{u_{i}\},\mathcal{V}^{\prime}\cup\{v_{i}\},\{f_{i}\})$ with causal mechanism $v_{i}=f_{i}(u_{i},\text{pa}_{\mathcal{V}^{\prime}}(v_{i}))$, where $u_{i}\in\mathcal{U}$ is an exogenous variable representing latent, uninterpretable information (e.g., noise), $v_{i}\in\mathcal{V}$ is an endogenous variable representing interpretable, symbolic information, and $f_{i}:U\times V\to V$ is a function describing the causal mechanism that determines the value of $v_{i}$ given its parents $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$. Such dependencies are captured by a graph $\mathcal{G}=(\mathcal{V}\cup\mathcal{V}^{\prime},\{(a,b)\mid a\in\mathcal{V}^{\prime},b\in\mathcal{V}\})$, representing all direct causal dependencies between endogenous variables.
Remark 3.4.
The structure of dissected Causal CEMs resembles the structure of CBMs as the prediction of each $v_{i}$ can be traced back to $v_{j}^{\prime},\ \forall j\neq i$. In this sense, endogenous copies in a Causal CEM play the role of “explaining variables”, akin to a CBM’s concepts. In contrast, endogenous variables play the role of “explained variables”, akin to a CBM’s tasks.
We can interpret the factors of a dissected Causal CEM as follows: $p(u\mid x)$ is the exogenous encoder, i.e., a deterministic distribution that is parametrised by a neural network $\zeta:X\to U$. In CBMs, this function represents the input encoder. The exogenous encoder $\zeta$ generates exogenous variables $u_{i}\in\mathcal{U}$ mapping raw input features $\mathbf{x}$ (e.g., an image’s pixels) to latent embeddings $\hat{\mathbf{u}}_{i}\in\mathbb{R}^{q},\ q\in\mathbb{N}$. In practice, this process mirrors the generation of context vectors in Concept Embedding Models [36]. First, the encoder $\psi:X\rightarrow H$ maps raw features to a latent code $\mathbf{h}\in H$. Then, a pair of neural networks $\{\phi_{i}^{+},\phi_{i}^{-}\}$ map the latent code into two different embeddings whose concatenation $[\phi_{i}^{+}(\mathbf{h}),\phi_{i}^{-}(\mathbf{h})]^{T}$ corresponds to the exogenous variable $\mathcal{U}_{i}$ of the $i$-th concept:
In contrast, $p(v^{\prime}\mid u)$ is the copies predictor. This is the product of $k$ independent Bernoulli distributions whose logits are parameterised by a neural network $s:U\to V$. In CBMs, the composition of the exogenous encoder and the concept predictor is often called concept encoder $g=\zeta\circ s$. In causal methods, this function represents a (supervised) causal feature learner [7]. The copies predictor $s$ generates endogenous copies $v_{i}^{\prime}\in\mathcal{V}$ from latent embeddings $\hat{\mathbf{u}}_{i}$. This is obtained using a neural network classifier $s:U\to V$ as a scoring function as in [36]:
Finally, $p(v_{i}\mid\text{pa}_{\mathcal{V}^{\prime}}(v_{i}),u_{i})$ is the endogenous predictor. This distribution is the product of $k$ independent Bernoulli distributions whose logits are parameterised by a neural network $f:V^{k}\times U\to V$. The input to this function $\text{pa}_{\mathcal{V}^{\prime}}(v_{i})$ (representing direct causal dependencies) is weighted by a learnable adjacency matrix $M\subseteq\mathbb{R}^{k\times k}$, where each learnable weight $m_{ij}$ models the strength of the dependency of $v_{i}$ from its parents $v_{j}^{\prime}$. In CBMs this function is called a task predictor [23] and intuitively represents the analogous of the structural equations that model causal mechanisms in SCMs [7]. The endogenous predictor generates the endogenous variables $\hat{v}_{i}\in\mathcal{V}$ by considering exogenous variables $\hat{\mathbf{u}}_{j}$ and copies $\hat{v}_{j}^{\prime}$ with $j\neq i$. First, the function $\omega:V\times U\to\mathbb{R}^{z}$ generates endogenous embeddings $\hat{\mathbf{v}}_{j}^{\prime}$ using exogenous variables $\hat{\mathbf{u}}_{j}$ and copies $\hat{v}_{j}^{\prime}$, following [36]. Then, all endogenous embeddings are weighted by the strength of the dependency $m_{ij}$ and aggregated using a deepset-like neural network $f_{i}:\mathbb{R}^{z\times k}\to[0,1]$ which maps endogenous embeddings to endogenous predictions:
(endogenous variables) $\quad\hat{v}_{i}=f_{i}\left(\{m_{ij}\hat{\mathbf{v}}_{j}^{\prime}\}_{j\in\{1,\dots,k\}}\right).$ (6)
In order to learn explicit structural equations, existing logic-based aggregation methods can be used from the concept literature [38, 33]. App. A describes in more detail their adaptation in Causal CEM.
3.2 Training Causal CEMs
[…]
Figure 2: Unfolded Causal CEM’s endogenous predictor.
Learning causal structures
@@ 3.2 Training Causal CEMs
Causal CEMs initialise the weights of causal dependencies $m_{ij}$ based on the conditional entropy between labels $v_{i}$ and $v_{j}$, as it can be used to represent asymmetric concept relationship (for other initialisation strategies, see App. A). These weights are then fine-tuned through an end-to-end learning process within the Causal CEM framework following common causal priors, where causal graphs are assumed to be sparse, directed, and acyclic (forming a Directed Acyclic Graph or DAG). We introduce a parameter $\gamma\in\mathbb{R}$ to eliminate less significant dependencies and a loss function, as described by Yang et al. [39], to enforce the sparsity and acyclicity of the causal graph, ensuring that the adjacency matrix $A$ effectively represents a DAG:
(initialization) $\quad m_{ij}=-\sum\nolimits_{a\in\{0,1\}}\sum\nolimits_{b\in\{0,1\}}p(v_{i}=a,v_{j}=b)\log p(v_{i}=a\mid v_{j}=b)$ (7)
@@
Remark 3.5 (Unfolding Causal CEMs with directed message passing).
Notice how learning a DAG together with Definition 3.2 allows to unfold a Causal CEM’s endogenous predictor applying a directed message passing on the associated structural causal model $\mathcal{M}$, ensuring that the values of endogenous variables are derived solely from the nodes that are their ancestors on the causal graph (see Figure 2). As a first step, we compute the exogenous variables for all nodes. We then predict the values of endogenous variables in root nodes in the learned DAG from their corresponding exogenous variables. Following this, we can generate the endogenous embeddings for root nodes and aggregate endogenous embeddings to compute the value of endogenous variables of each child node. We repeat this process until all leaf nodes of the graph are reached. We can obtain this by replacing in Eq. 5 the endogenous copies $\hat{v}_{j}^{\prime}$ with the parents of the endogenous variable $v_{i}$:
Note that this causal unrolling guarantees two key properties (see Figure 3(a)): (1) modifying the value of a cause (parent node) will impact the effect (child node) in our model, (2) conversely, intervening upon an effect does not alter the cause. This is because, in our model, information flows sequentially following the graph, mirroring the fundamental nature of causal effects. Consequently, this layer not only facilitates the computation of task predictions but also enables the exploration of causal relationships through do-interventions and counterfactual analysis.
@@
We can now state the general learning objective for Causal CEMs. Given (1) a set of entities represented by their feature vectors $\mathbf{x}\in\mathcal{X}\subseteq X$ (i.e. the input) and (2) a set of annotations for each exogenous variable $v\in\mathcal{V}\subseteq V$ (i.e. the labels), we wish to find functions $\zeta$, $s$, $f$, together with the adjacency matrix $A$, that maximise the log-likelihood of $v,v^{\prime}$, while observing $x$ (or equivalently $u=\zeta(x)$):
@@
where $\lambda_{1,2}$ are hyperparameters balancing optimisation objectives. Notice that we use the $do$-operator replacing $\hat{v}_{j}^{\prime}$ with labels $v_{j}$ to minimise leakage and provide better gradients to the endogenous predictor $f$. This enables Causal CEMs to be aware of $do$-operations during training, thus making the model effective in responding to $do$-interventions once deployed.
4 Causal reasoning and verification with Causal CEM
As traditional CBMs, Causal CEMs enable ground-truth intervention that allow domain experts to fix mispredicted concept labels at test time. This notion of interventions must not be confused with the notion of intervention usually discussed in causality literature and modelled by Pearl’s do-operator. For the latter, here we use the term do-intervention. However, in the proposed architecture, the ground-truth interventions have a potentially higher impact as each intervention has a downstream effect on all descendant nodes in the causal graph. In contrast, a concept intervention in a traditional CBM has only an effect on 1-hop nodes representing the downstream task. Moreover, Causal CEMs enable sound causal inference and verification via do-interventions and counterfactual analysis.
Ground-truth interventions Causal CEMs support “ground-truth interventions” (see Figure 3(b)). Ground-truth interventions are one of the core motivations behind CBMs [23]. Through ground-truth interventions, concept bottleneck models allow experts to improve a CBM’s task performance by rectifying mispredicted concepts at test time, thus significantly improving task performance within a human-in-the-loop setting. In Causal CEMs, however, ground-truth interventions have a potential impact on all endogenous variables descendant of an intervened node, which may include not only nodes corresponding to downstream tasks but also nodes corresponding to a CBM’s intermediate concepts. This enables a single concept ground-truth intervention to potentially improve the prediction of intermediate concepts as well as downstream tasks.
@@ 4 Causal reasoning and verification with Causal CEM
Causal CEMs can answer interventional and counterfactual queries related to the model’s decision-making process using the do-operator on the unfolded SCM.
@@
Do-interventions enable manipulation of a Causal CEM’s decision-making process by changing the value of a specific endogenous variable and observing how it affects other variables’ distributions (see Figure 3(c)). In Causal CEMs, the effect of the $do$-intervention is analysed through the interventional distribution, denoted as $p(v_{i}|do(v_{j}=\kappa),\text{pa}(v_{i}))$, which describes the distribution of the outcome variable $v_{i}$ after the intervention $do(v_{j}=\kappa)$ has been performed. In particular, in Causal CEMs, the do-operation fixes the value of the intervened variable $v_{i}$ to a fixed constant $\kappa\in\{0,1\}$ and removes all causal dependencies from parent variables by zeroing all values of the $i$-th column of the adjacency matrix:
@@
Causal CEMs also enable to answer counterfactual queries such as “What would the value of the $i$-th variable have been, had the $j$-th variable been $\kappa$, given that we observed $v_{i}$ and $v_{j}$?”. Answering these queries involves three steps [27]: 1) Abduction: Infer a realisation of exogenous variables that is consistent with the observed $v_{i}$ and $v_{j}$ in the actual causal model. 2) Action: Modify the architecture of the causal CEM $\mathcal{M}$ into $\mathcal{M}_{v_{j}=\kappa}$ by replacing the structural equation for $v_{j}$ with $\kappa$, to simulate the intervention. 3) Prediction: Compute the value of $v_{i}$ in the modified model $\mathcal{M}_{\kappa}$, representing the counterfactual outcome:
(abduction) $\quad\hat{\mathbf{u}}_{i}=\zeta(\mathbf{x})$ (12)
@@
(prediction) $\quad\hat{v}_{i}=f_{i}\left(\{a_{in}\omega(\hat{v}_{n},\hat{\mathbf{u}}_{n})\}\right),\quad\forall i,n\in\mathcal{V}$ (14)
@@
Model verification and blocking Causal analysis enables the verification of properties of Causal CEMs before deployment. For instance, using only the learnt causal graph, it is possible to prove that an endogenous variable $v_{i}$ is independent of the variable $v_{j}$ by verifying that $v_{j}$ is not among the ancestors of $v_{i}$. Another form of formal verification, which we call “blocking”, employs the $do$-intervention (see Figure 3(d)). Blocking allows one to formally verify the independence of a pair of variables given a sequence of $do$-operations. Given a pair of variables $v_{j}$ and $v_{i}$ such that $v_{j}$ is an ancestor of $v_{i}$, we perform a blocking verification as follows: 1) Block: perform a do-intervention on all child nodes of $v_{j}$, 2) Verify: perform a do-operation on $v_{j}$ itself and observe the impact on $v_{i}$. We can easily verify that the first step makes $v_{j}$ and $v_{i}$ completely independent by observing that the do-operation on $v_{j}$ no longer alters the distribution of $v_{j}$.
@@ 5 Experiments
  • Concept-based performance and interpretability: Can Causal CEMs match the generalisation performance of equivalent black-box models and existing CBMs? Can Causal CEMs enable more effective ground-truth interventions w.r.t. existing CBMs?
  • Causal Interpretability: Are Causal CEMs causally interpretable? Can Causal CEMs effectively block the causal effect of two causally related endogenous variables?
To answer these questions, we use three datasets: (i) Checkmark, a synthetic dataset composed of four endogenous variables; (ii) dSprites, where endogenous variables correspond to object types together with their position, colour, and shape; and (iii) CelebA, a facial recognition dataset where endogenous variables represent facial attributes. Using these datasets, we compare the proposed approach with a black box baseline and state-of-the-art concept-based architectures: Concept Bottleneck Models (CBM[23], and Concept Embedding Models (CEM[36]. We also compare a version of the proposed method (Causal CEM) where the causal graph is learnt end-to-end w.r.t. with a version (Causal CEM+CD) where the causal graph is extracted from ground-truth labels using a causal discovery algorithm [40]. We provide further details on our experimental setup and baselines in App. B.
A comprehensive set of experiments is detailed in App. C, where the experiments presented in this section for a subset of the datasets are extended to include all datasets.
@@ 5.1 Key findings
checkmark dsprites celeba
Black box $90.15{\scriptstyle\pm 1.30}$ $99.53{\scriptstyle\pm 0.05}$ $79.55{\scriptstyle\pm 0.14}$
CBM $90.34{\scriptstyle\pm 0.55}$ $99.55{\scriptstyle\pm 0.07}$ $79.00{\scriptstyle\pm 0.18}$
CEM $89.09{\scriptstyle\pm 1.98}$ $99.48{\scriptstyle\pm 0.07}$ $79.17{\scriptstyle\pm 0.26}$
Causal CEM+CD $89.43{\scriptstyle\pm 0.93}$ $99.40{\scriptstyle\pm 0.15}$ $78.42{\scriptstyle\pm 0.42}$
Causal CEM $88.24{\scriptstyle\pm 1.30}$ $99.44{\scriptstyle\pm 0.11}$ $78.23{\scriptstyle\pm 0.45}$
Causal CEMs match the performance of causally opaque models. (Table 1)
Causal CEMs demonstrate robust generalisation across all datasets, yielding a predictive performance close to that of black-box architectures with an equivalent capacity. Causal CEMs using a pre-trained causal graph (Causal CEM+CD) tend to have slightly better label accuracy with respect to Causal CEMs where the causal graph is learned end-to-end (Causal CEM). Causal CEMs’s low variance suggests a consistent robustness on weight initializations over multiple training runs.
Figure 4: Impact of ground-truth interventions on non-intervened nodes ($\uparrow$).
@@
Ground-truth interventions on Causal CEMs improve both concept and task accuracy as opposed to CBMs (Figure 4)
@@
In Causal CEM, the causal graph induces a natural strategy for ground-truth interventions. Indeed, the causal graph narrows down the set of variables to intervene upon: for any given node, we can just fix mispredicted labels of the node’s ancestors as intervening on other nodes will not have any impact. This property significantly decreases the required number of interventions to achieve a desired outcome (e.g., to increase a downstream task accuracy), as shown in App. C. Another advantage of Causal CEM consists in the hierarchical nature of inference which allows ground-truth interventions to impact all endogenous variables descendant of an intervened node. In particular, ground-truth interventions may affect not only nodes corresponding to downstream tasks (as in CBM and CEM), but also nodes corresponding to a CBM’s intermediate concepts. We experimentally verify this property and its impact by calculating—for nodes that were not intervened upon (including both concepts and tasks)—the change in accuracy before and after ground-truth interventions were applied on their ancestors. Our results (Figure 4) show that Causal CEM improves nodes accuracy by $\sim 15$ percentage points after only 7 ground-truth interventions on CelebA, while CBM and CEM node accuracy remains almost unchanged. Causal CEM’s advantage increases with the number of concepts and connections, as a single intervention can impact a higher number of nodes in the causal graph. CBM and CEM, instead, achieve a similar performance only after intervening on all concepts as their architecture assumes all concepts to be mutually independent.
Causal CEMs’ endogenous predictors are causally interpretable (Figure 5)
In Causal CEM the decision-making process is causally interpretable and can be analysed by visualising the learnt causal graph and structural equations as in a structural causal model, as shown in Figure 5 for CelebA. The image shows how Causal CEM exploited known biases in CelebA to infer facial attributes. For instance, Causal CEM predicts the attribute “wearing lipstick” directly from the attribute “attractive” and indirectly from attributes such as “smiling” or “high cheek”, all attributes that are known to be strongly correlated with each other in CelebA [41, 42].
Figure 5: Portion of the learnt causal graph and structural equations in CelebA. A node’s colour in the causal graph is proportional to the probability of necessity and sufficiency w.r.t. the node $v_{5}$.

### v4

@@ 3.2 Causal transparency in Concept Graph Models
Figure 2: $p(v_{1}\mid v_{k}^{\prime},u_{1};\theta_{f})$ is equivalent to $p(v_{1}\mid v_{k},u_{1};\theta_{f})$ as they both represent the same query on the same conditional probability table.
The above theorem shows that CGM’s bipartite graph $\mathcal{G}^{\prime}$ between nodes $\mathcal{V}^{\prime}$ and $\mathcal{V}$ can efficiently represent arbitrary long cause-effect chains among interpretable variables (example in Fig. 2). While during training CGMs exploit the efficient bipartite representation, at test time CGMs can be unfolded following the graph as described in Theorem 3.2 (Sec. 3.3 describes how to learn sparse and acyclic dependencies to make the test time unfolding more efficient and interpretable; we provide more details in App. H.6). This property makes CGMs causally transparent by design, as they can make class predictions through cause-effect chains of endogenous variables, allowing humans to trace each prediction back to its underlying causes within an interpretable parental subgraph. Notice how—unlike CBMs—Causal CGMs do not require to know in advance cause (concept)-effect (task) relationships, nor require that all causes (concepts) are independent from each other, as cause-effect dependencies among interpretable variables are not given, but learnt during training.
@@
Causal CGMs are designed to address the problem of causal opacity, not causal discovery. Causal opacity—the problem of understanding a model’s decision-making process— neither implies nor requires causal discovery—the problem of identifying the rules governing the world or the data-generating mechanism. Indeed, the causal structure of a model’s decision-making process does not necessarily mirror the causal structure of the world (Boge, 2022). We report in appendices extended discussions on CGM’s relations with causal discovery (App. J) and causal representation learning (App. K), and on CGM’s identifiability (App. L).
3.3 Training Causal CGMs
Learning sparse, directed, and acyclic structures
@@ 3.3 Training Causal CGMs
In order to construct simple explanations, we can make Causal CGMs’ graphs sparse and acyclic. To this end, we parametrize the bipartite graph $\mathcal{G}^{\prime}$ with an adjacency matrix of learnable weights $M\in\mathbb{R}^{k\times k}$. We initialise the weights $m_{ij}\in M$ based on the conditional entropy between labels $v_{i}$ and $v_{j}$ in order to account for asymmetric concept dependencies (for other initialisation strategies, see App. F). These weights are then fine-tuned through an end-to-end learning process within the Causal CGM framework following common causal priors, where causal graphs are assumed to be sparse, directed, and acyclic (forming a Directed Acyclic Graph or DAG). We introduce a parameter $\gamma\in\mathbb{R}$ to eliminate less significant dependencies and a loss function, as described by Yang et al. (2020), to enforce the sparsity and acyclicity of the causal graph, ensuring that the adjacency matrix $A$ effectively represents a DAG:
(initialisation) $\quad a_{ij}=-\sum\nolimits_{b\in\{0,1\}}\sum\nolimits_{c\in\{0,1\}}p(v_{i}=b,v_{j}=c)\log p(v_{i}=b\mid v_{j}=c)$ (2)
@@
We can now state the general learning objective for Causal CGMs. Given (1) a set of entities represented by their feature vectors $\mathbf{x}\in\mathcal{X}\subseteq X$ (i.e. the input) and (2) a set of annotations for each exogenous variable $v\in\mathcal{V}\subseteq V$ (i.e. the labels), we wish to find functions $\zeta$, $s$, $f$, together with the adjacency matrix $A$, that maximise the log-likelihood of $v,v^{\prime}$, while observing $x$ (or equivalently $u=\zeta(x)$):
@@
where $\lambda_{1,2}$ are hyperparameters balancing optimisation objectives (related ablation studies can be found in App. H.3). Notice that we use the $do$-operator replacing $\hat{v}_{j}^{\prime}$ with labels $v_{j}$ to minimise leakage and provide better gradients to the endogenous predictor $f$. This enables Causal CGMs to be aware of $do$-operations during training, thus making the model effective in responding to $do$-interventions once deployed.
3.4 Causal reasoning and verification with Causal CGM
CGMs support two kinds of interventions: ground-truth interventions, as in CBMs, and do-interventions, as in causal models. In what follows, we first highlight how CGM’s design makes ground-truth interventions more effective and then we describe how do-interventions can be applied in Causal CGMs to verify properties of the model’s decision-making process.
Ground-truth interventions
Causal CGMs support “ground-truth interventions” (see Figure 3(b)). Ground-truth interventions are one of the core motivations behind CBMs (Koh et al., 2020). Through ground-truth interventions, concept bottleneck models allow experts to improve a CBM’s task performance by rectifying mispredicted concepts at test time, thus significantly improving task performance within a human-in-the-loop setting. In Causal CGMs, however, ground-truth interventions have a potential impact on all endogenous variables descendant of an intervened node, which may include not only nodes corresponding to downstream tasks but also nodes corresponding to a CBM’s intermediate concepts. This enables a single concept ground-truth intervention to potentially improve the prediction of intermediate concepts as well as downstream tasks.
@@ 3.4 Causal reasoning and verification with Causal CGM
Causal CGMs can answer interventional and counterfactual queries related to the model’s decision-making process using the do-operator on the unfolded SCM.
@@
Do-interventions enable manipulation of a Causal CGM’s decision-making process by changing the value of a specific endogenous variable and observing how it affects other variables’ distributions (see Figure 3(c)). In Causal CGMs, the effect of the $do$-intervention is analysed through the interventional distribution, denoted as $p(v_{i}|do(v_{j}=\kappa),\text{pa}(v_{i}))$, which describes the distribution of the outcome variable $v_{i}$ after the intervention $do(v_{j}=\kappa)$ has been performed. In particular, in Causal CGMs, the do-operation fixes the value of the intervened variable $v_{i}$ to a fixed constant $\kappa\in\{0,1\}$ and removes all causal dependencies from parent variables by zeroing all values of the $i$-th column of the adjacency matrix:
@@
Causal CGMs also enable to answer counterfactual queries such as “What would the value of the $i$-th variable have been, had the $j$-th variable been $\kappa$, given that we observed $v_{i}$ and $v_{j}$?” (extended motivation in App I). Answering these queries involves three steps (Pearl, 1995): 1) Abduction: Infer a realisation of exogenous variables that is consistent with the observed $v_{i}$ and $v_{j}$ in the actual causal model. 2) Action: Modify the architecture of the Causal CGM $\mathcal{M}$ into $\mathcal{M}_{v_{j}=\kappa}$ by replacing the structural equation for $v_{j}$ with $\kappa$, to simulate the intervention. 3) Prediction: Compute the value of $v_{i}$ in the modified model $\mathcal{M}_{\kappa}$, representing the counterfactual outcome:
(abduction) $\quad\hat{u}_{j}=p(u_{j}\mid\mathbf{x};\theta_{\zeta})$ (6)
@@
(prediction) $\quad\hat{v}_{i}=p(\hat{v_{i}}\mid\text{pa}_{\mathcal{V}}(v_{i}),u_{i};\theta_{f})\ \forall i\neq j$ (8)
@@
Model verification and blocking Causal analysis enables the verification of properties of Causal CGMs before deployment. For instance, using only the learnt causal graph, it is possible to prove that an endogenous variable $v_{i}$ is independent of the variable $v_{j}$ by verifying that $v_{j}$ is not among the ancestors of $v_{i}$. Another form of formal verification, which we call “blocking”, employs the $do$-intervention (see Figure 3(d)). Blocking allows one to formally verify the independence of a pair of variables given a sequence of $do$-operations. Given a pair of variables $v_{j}$ and $v_{i}$ such that $v_{j}$ is an ancestor of $v_{i}$, we perform a blocking verification as follows: 1) Block: perform a do-intervention on all child nodes of $v_{j}$, 2) Verify: perform a do-operation on $v_{j}$ itself and observe the impact on $v_{i}$. We can easily verify that the first step makes $v_{j}$ and $v_{i}$ completely independent by observing that the do-operation on $v_{j}$ no longer alters the distribution of $v_{j}$.
@@ 4 Experiments
  • Concept-based performance and interpretability: Can Causal CGMs match the generalisation performance of equivalent black-box models and existing CBMs? Can Causal CGMs enable more effective ground-truth interventions w.r.t. existing CBMs?
  • Causal Interpretability: Are Causal CGMs causally interpretable? Can Causal CGMs effectively block the causal effect of two causally related endogenous variables?
To answer these questions, we use four datasets: (i) Checkmark, a synthetic dataset composed of four endogenous variables; (ii) dSprites, where endogenous variables correspond to object types together with their position, colour, and shape; (iii) CelebA, a facial recognition dataset where endogenous variables represent facial attributes; (iv) CIFAR10, an animal classification dataset where the endogenous variables are extracted automatically following Oikarinen et al. (2023). Using these datasets, we compare the proposed approach with a black box baseline and state-of-the-art concept-based architectures: Concept Bottleneck Models (CBM(Koh et al., 2020), and Concept Embedding Models (CEM(Espinosa Zarlenga et al., 2022). We also compare a version of the proposed method (Causal CGM) where the causal graph is learnt end-to-end w.r.t. with a version (Causal CGM+CD) where the causal graph is extracted from ground-truth labels using a causal discovery algorithm (Lam et al., 2022) and injected into the model fixing the matrix $A$. Due to the prohibitive computational complexity of the causal discovery method, this was not feasible for the CIFAR-10 dataset We provide further details on our experimental setup and baselines in App. G.
A comprehensive set of experiments is detailed in App. H, where the experiments presented in this section for a subset of the datasets are extended to include all datasets. We also report an analysis of Causal CGMs’ scalability and computational complexity in App. H.5.
We remark that, following the CBM literature, the goal of Causal CGMs is not to increase performance but interpretability. Therefore, the experiments aim to show that our Causal CGMs’ classification performance is competitive against existing black-box models and state-of-the-art concept-based approaches (e.g., within one percentage point) while enabling better visibility and manipulation of its reasoning process.
@@ 4.1 Key findings
Causal CGMs match the performance of causally opaque models. (Table 1)
Causal CGMs demonstrate robust generalisation across all datasets, yielding a predictive performance close to that of black-box architectures with an equivalent capacity. Causal CGMs using a pre-trained causal graph (Causal CGM+CD) tend to have slightly better label accuracy with respect to Causal CGMs where the causal graph is learned end-to-end (Causal CGM). Causal CGMs’s low variance suggests a consistent robustness with respect to weight initialisations over multiple training runs. Thanks to concept embeddings, concept incompleteness settings do not have a strong impact on Causal CGMs’ performance (results in App. H.4).
@@
checkmark dsprites celeba cifar10 semantic transparency causal transparency
Black box $90.15{\scriptstyle\pm 1.30}$ $99.53{\scriptstyle\pm 0.05}$ $79.55{\scriptstyle\pm 0.14}$ $94.85{\scriptstyle\pm 0.03}$
CBM $90.34{\scriptstyle\pm 0.55}$ $99.55{\scriptstyle\pm 0.07}$ $79.00{\scriptstyle\pm 0.18}$ $92.17{\scriptstyle\pm 0.11}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
CEM $89.09{\scriptstyle\pm 1.98}$ $99.48{\scriptstyle\pm 0.07}$ $79.17{\scriptstyle\pm 0.26}$ $92.04{\scriptstyle\pm 0.06}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Causal CGM+CD $89.43{\scriptstyle\pm 0.93}$ $99.40{\scriptstyle\pm 0.15}$ $78.42{\scriptstyle\pm 0.42}$ N/A ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Causal CGM $88.24{\scriptstyle\pm 1.30}$ $99.44{\scriptstyle\pm 0.11}$ $78.23{\scriptstyle\pm 0.45}$ $93.32{\scriptstyle\pm 0.03}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
@@
Ground-truth interventions on Causal CGMs improve both concept and task accuracy as opposed to CBMs (Figures 5, 5)
@@
In Causal CGM, the causal graph induces a natural strategy for ground-truth interventions. Indeed, the causal graph narrows down the set of variables to intervene upon: for any given node, we can just fix mispredicted labels of the node’s ancestors as intervening on other nodes will not have any impact. This property significantly decreases the required number of interventions to achieve a desired outcome (e.g., to increase a downstream task accuracy), as shown in App. H. Another advantage of Causal CGM consists in the hierarchical nature of inference which allows ground-truth interventions to impact all endogenous variables descendant of an intervened node. In particular, ground-truth interventions may affect not only nodes corresponding to downstream tasks (as in CBM and CEM), but also nodes corresponding to a CBM’s intermediate concepts. We experimentally verify this property and its impact by calculating—for nodes that were not intervened upon (including both concepts and tasks)—the change in accuracy before and after ground-truth interventions were applied on their ancestors. Our results (Figure 5 and App. H.1) show that Causal CGM improves nodes accuracy by $\sim 15$ percentage points after only 7 ground-truth interventions on CelebA. CBM and CEM, instead, achieve a similar performance only after intervening on all concepts as their architecture assumes all concepts to be mutually independent. Focusing solely on concept-level analysis, Causal CGM enhances concept accuracy by up to $\sim 25$ percentage points, compared to CBM and CEM, where a concept intervention does not impact the accuracy of other concepts (Figure 5). Causal CGM’s advantage increases with the number of concepts and connections, as a single intervention can impact a higher number of nodes in the causal graph.
Figure 4: Impact of ground-truth interventions on non-intervened nodes ($\uparrow$). Intervention on Causal CGM improves both concept and task accuracy.
Figure 5: Impact of ground-truth interventions on concept nodes ($\uparrow$). Intervention on CBM and CEM do not improve concept accuracy unlike Causal CGM.
Causal CGMs’ endogenous predictors are causally interpretable (Figures 6(a), 6(b))